If the consumption function is C = $400 billion + 0.8Y Note: some values are in trillions and others are in billions (a) What is the MPC Instructions: Enter your response rounded to one decimal place. Instructions: Enter your response rounded to the nearest whole number. (b) How large is autonomous C? billion (c) How much do consumers spend with incomes of $4 trillion? billion (d) How much do they save? billion
media%2Fe85%2Fe859802a-df5f-494c-be1e-8c
Answer #1

Consumption function = Autonomus C + MPC ( Y)

Here, Consumption function = $400 billion + 0.8Y

a) MPC = 0.8

b) Autonomous Consumption = $400 billion

c) Consumer spend with income of $4 trillion ( $4 trillion = $4000 billion)

   Spending = Consumption as whatever money is spent is used for consumption only

   Putting value of Y in Consumption function

   Consumption function = $400 billion + 0.8 ($4000 billion)

                                    = $400 billion + $3200 billion

                                    = $3600 billion

d) First deriving saving function

    Saving function = (-Autonomous Consumption) + MPS ( Y)

       MPS = 1 - MPC = 1 - 0.8 = 0.2

   Saving function = -$400 billion + 0.2 ($4000 billion)

                          = -$400 billion + $800 billion

                          = $400 billion